OpenStudy (mendicant_bias):

Does taking the derivative of a system of equations affect the range of those equations? I've never had to do this before, so just asking. (Example below.)

OpenStudy (mendicant_bias):

\[g(x) = \left\{ x^3, -2\le x \le 0 \right\}\]\[g(x) =\left\{ x^2, 0<x \le 2 \right\}\](Improper use of brackets, I know, but I can't write it the "proper" way yet with Openstudy.) \[g'(x) = 3x ^{2}\]\[g'(x) = 2x\]Are the ranges of these at all affected?

OpenStudy (anonymous):

the ranges are not effected at all, what the original equation is stating is that it is restricting them to their respective domains.

OpenStudy (mendicant_bias):

Just making sure. I know what the original equation is stating, but I don't want to assume that continually holds true. Thanks.